OpenStudy (anonymous):

simplify 4x^2 -32x + 48 / 3x^2 -17x -6

OpenStudy (anonymous):

Factorize first numerator and denominator..

OpenStudy (anonymous):

Take 4 common from the numeration and then factorize it, this will be simple to factorize..

OpenStudy (anonymous):

for the numerator i got 4(x-6)(x-2) but how are you suppose to simplify the denominator when 17 cant be simplified?

OpenStudy (anonymous):

Denominator is: \[= (3x +1)(x-6)\]

OpenStudy (anonymous):

[4(x-6)(x-2)/ (x-6)(3x+1)]=0

OpenStudy (anonymous):

how did you guys get the denominator?

OpenStudy (anonymous):

Cancel what you can cancel and then tell me what you are left with??

OpenStudy (anonymous):

It is simple: See, when: \[a \ne 1\] Then we make the factors of ac.. that is : 3*(-6) = -18 Getting??

OpenStudy (anonymous):

Ok I see what your saying

OpenStudy (anonymous):

In numerator after taking 4 common the coefficient of x became 1 but in denominator coefficient of x is not 1 but 3.. In this case we multiply coefficient of x with constant term that is c.. So, 3*(-6) = 18 Factors of 18 can be 18 and 1.. Can you proceed further??

OpenStudy (anonymous):

yes thank you ;D

OpenStudy (anonymous):

Welcome dear..